OpenStudy (anonymous):

Anyone up for a challenge? Find the area bounded by these three curves: y=4-4x; y=2xcos(x^2); y=sin(x)cos^2(x) in the first quadrant. Use the calculator to get the intersection points...I just can't seem to get the same answer twice! Algebra issues...

OpenStudy (anonymous):

It's the limits of integration, you know plugging everything in that's getting me. Let me know the answers you get.

OpenStudy (anonymous):

from the picture it looks like the left hand endpoint is where \[2x\cos^2(x)=\sin(x)\cos^2(x)\] i.e. \[x=0\]

OpenStudy (anonymous):

Yes, I agree.

OpenStudy (anonymous):

The endpoints I get for the other two intersections are .6928 and 1.237

OpenStudy (anonymous):

( I just plugged the functions into my graphing calculator and evaluated the intersections where it seemed that the graphs overlapped.)

OpenStudy (anonymous):

then next one is where \[4-4x=2x\cos^2(x)\] wolfram gives me \[x= 0.708523315642791...\]

OpenStudy (anonymous):

Really! I should have used wolfram. Let me try it again with new numbers.

OpenStudy (anonymous):

still have to find where \[4-4x=\sin(x)\cos^2(x)\]

OpenStudy (anonymous):

We can do that the same way--wolfram, right?

OpenStudy (anonymous):

yeah i got \[0.928113\] http://www.wolframalpha.com/input/?i=4-4x%3Dsin%28x%29cos^2%28x%29

OpenStudy (anonymous):

Hmm. The first one is actually 2xcos(x^2)

OpenStudy (anonymous):

It's not 2xcos^2(x). So, 4-4x=2xcos(x^2) = .6928

OpenStudy (anonymous):

So, 0, .6928, .9281

OpenStudy (anonymous):

You found, accidentally 4-4x=2xcos^x(x) and got .7085, I think.

OpenStudy (anonymous):

oops i tried it again and got this http://www.wolframalpha.com/input/?i=2xcos^2%28x%29+%3D+4-4x

OpenStudy (anonymous):

No no! It's 2xcos(x^2). Not cosine squared of x, but cosine of x squared. ><

OpenStudy (anonymous):

oh lord

OpenStudy (anonymous):

.6828 it is

OpenStudy (anonymous):

It's okay :P i just found out how to evaluate the definite integral on my calculator, so I might be able to get this now! Doing it by hand was TOO much.

OpenStudy (anonymous):

But if you'd do it too so we can check answers...

OpenStudy (anonymous):

forget doing it by hand

OpenStudy (anonymous):

That'd be awesome.

OpenStudy (anonymous):

OH I KNOW. It was just awful.

OpenStudy (anonymous):

now we just have to figure out what to integrate

OpenStudy (anonymous):

mhm. I split it up. So from 0 to .6928, you do 2xcos(x^2)-(4-4x)?

OpenStudy (anonymous):

no, not from the picture looks like you should do this http://www.wolframalpha.com/input/?i=integral+0+to+0.692751+%282xcos^2%28x%29-sin%28x%29cos^2%28x%29%29

OpenStudy (anonymous):

No. no. 2xcos(x^2)-sinxcos^2(x)

OpenStudy (anonymous):

Yeah, what you said :P

OpenStudy (anonymous):

whew

OpenStudy (anonymous):

over that interval \[2x\cos(x^2)>\sin(x)\cos^2(x)\]

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

HEY SATELLITE! We got it right! It was an option! Nice! Thank you so much :)

OpenStudy (anonymous):

wait!!

OpenStudy (anonymous):

we still have to do the second part

OpenStudy (anonymous):

Oh, I added the other side already

OpenStudy (anonymous):

.378 something is the anwer.

OpenStudy (anonymous):

Don't worry :) I got it. thanks for the suggestion of using wolfram, though. It's much easier than typing all of this in the calculator.

OpenStudy (anonymous):

yw but that is not what i get

OpenStudy (anonymous):

i got .195.. for the first part, and .098 for the second part, but i could be wrong

OpenStudy (anonymous):

If you add both together? The first section = approximately .208 and the second part = approximately .0983. You add them together to equal .378. Which was the correct answer.

OpenStudy (anonymous):

but if you got the right answer great!

OpenStudy (anonymous):

We differ on the first part.

OpenStudy (anonymous):

Might have to do with rounding. I dunno. But thanks so much for the help :)

OpenStudy (anonymous):

yw